Need help with knowing how to do the math
Please use the following information to answer the next three questions: A currency speculator expects the spot rate of Euros to change from $1.00 to $0.80 in six months. Assume the speculator has access to credit lines of $10,000,000 in the US and EUR 10,000.000 in Europe. The annual borrowing& lending rates are 6 percent in US and 8 percent in Europe. (30) In order for the speculator to take advantage from the expected spot rate change in Euros. A. borrow in US $ and invest in Euros B. borrow in Euros and invest in US $ 31) If his forecast turns out be to true, at the end of the six month period, the speculators expected profit (in A. $2,475,000 US $) will be: B. $2,450,000 C. $1,980,000 D. $1,960,000
Answer #1

30. B Borrow in Euros and invest in US $

The rule is that if the interest rate differential is greater than the premium or discount, place the money in the currency that has a higher rate of interest or vice-versa.

Interest rate differential= 8-6= 2%

Forward premia (annualised)= (Forward rate- Spot Rate) *100*12/ Spot rate*6

= ($0.80- $1)*100*12/1*6= (40%)

31. C. $1,980,000

Step 1.
Borrow EUR 10,000,000 in Europe @8% for 6 months
Interest
10,000,000
400000
EURO10,400,000
Step 2
Convert 10,000,000 into $
10000000*$1
$10000000
Step 3
Invest $10,000,000 @ 6% for 6 months
$
10,000,000
300000
$10,300,000
Step 4
Convert 10,300,000$ into Euro
$ 10,300,000*1.25
EURO12875000
Step 5
Gain
12875000- 10,400,000
EURO2,475,000
Euro 2475,000 * 0.80
$
1980000